OpenStudy (blake3299):

what is x-2/7=3 A.x = –3 2/7 B.x = –2 5/7 C.x = 2 5/7 D.x = 3 2/7

OpenStudy (anonymous):

Something is missing..

OpenStudy (sparrow2):

definetelle :D

OpenStudy (blake3299):

it suppose to be x-2/7=3

OpenStudy (sparrow2):

this?

OpenStudy (anonymous):

Add 2/7 both the sides to find x..

OpenStudy (blake3299):

the 2/7 is a fraction

OpenStudy (anonymous):

Oh, no no, you can't do that..

OpenStudy (anonymous):

\[\frac{x - 2}{7} = 3\] Is this okay?

OpenStudy (blake3299):

x is three right

OpenStudy (anonymous):

No...

OpenStudy (blake3299):

2 or -2

OpenStudy (anonymous):

Firstly ask your question properly, I have guessed your question, have I guessed it correctly??

OpenStudy (anonymous):

\[x - \frac{2}{7} = 3 \quad Or \quad \frac{x-2}{7} = 3\] Which one?

OpenStudy (blake3299):

the one on the left

OpenStudy (anonymous):

Okay, good..

OpenStudy (anonymous):

None of the options match your question then.. Check your question properly..

OpenStudy (blake3299):

i say it is 2

OpenStudy (anonymous):

I forget that You are GOD.. You can say anything.. :P

OpenStudy (anonymous):

Just add 2/7 to 3, and check if you get 2 or not..

OpenStudy (blake3299):

i say a

OpenStudy (blake3299):

i ment d

OpenStudy (anonymous):

D is good.. :)

OpenStudy (anonymous):

Sorry, my net was slow..